OpenStudy (ℐℵk℣ⱱøƴȡ◆):

GCF for the list of terms x^3, x^6, x^7

OpenStudy (dayakar):



OpenStudy (dayakar):

in above three terms which are common

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

3?

OpenStudy (dayakar):

no x * x * x is common

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

ohhhh x^3?

OpenStudy (dayakar):

G C F = x * x * x = x^3

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

ohh thank youuu